OpenStudy (anonymous):

The dollar value v(t) of a certain car model that is t years old is given by the following exponential function. V(t) = 29,900(0.80^t) Find the initial value of the car and the value after 12 years. Round your answers to the nearest dollar as necessary. Initial value: Value after 12 years:

OpenStudy (er.mohd.amir):

for initial valve put t=0 in eq and for 12 years put t=12 in eq.

OpenStudy (anonymous):

can you walk me through solving? I am having some trouble with this

OpenStudy (er.mohd.amir):

What is problem?

OpenStudy (anonymous):

I am having trouble solving it

OpenStudy (er.mohd.amir):

for initial value t=0 ,v(0)=29,900(0.80^0) and a^0=1 where a is any numbers use this

OpenStudy (mathmale):

V(t) = 29,900(0.80^t) is an EXPONENTIAL FUNCTION. 0.80 is the BASE of this exponential function, and t is the EXPONENT or POWER to which this base is raised. Either using your calculator or by and, evaluate 0.80^0, 0.80^2, 0.80^3, and so on. Now put your exponential function to use: If you substitute 12 for t, V(t) = 29,900(0.80^t) will give you the (depreciated) value of the car after 12 years. Please show your work. Thanks.

OpenStudy (mathmale):

What kind of calculator have you?

OpenStudy (anonymous):

I have a scientific calculator

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

Have you done exponentiation on your calculator before? Try 2^3. Answer should be 8. Try 5^2. Answer should be 25.

OpenStudy (anonymous):

Yes it works!

OpenStudy (mathmale):

Then evaluate 0.8^0, 0.8^1, 0.8^2, and so on.

OpenStudy (mathmale):

What is 0.8^2?

OpenStudy (anonymous):

0.64

OpenStudy (mathmale):

Right! So, you want to calculate the value of the car after 12 years. The pertinent function is V(t) = 29,900(0.80^t). Find V(12). To do this, calculate 0.8^12 first, and only then multiply that result by $29,900. V(12)=?

OpenStudy (mathmale):

this follows "order of operations" rules.

OpenStudy (anonymous):

So the answer would be 1,992.864825?

OpenStudy (anonymous):

And if I round that to the nearest dollar it would be 1,993?

OpenStudy (er.mohd.amir):

wrong calculation

OpenStudy (anonymous):

No sorry it would be 2054.712354. I put 29,000 on accident

OpenStudy (anonymous):

So the answer would be 2055?

OpenStudy (mathmale):

2055 what?

OpenStudy (anonymous):

$2,055 is the value after 12 years

OpenStudy (mathmale):

Cool. Perfect! Nice work! Happy New Year!

OpenStudy (mathmale):

Have to get off the 'Net now, but hope to be able to work with you again.

OpenStudy (anonymous):

Ok thanks!

OpenStudy (mathmale):

Sure! Bye.